# Quantitative Uniqueness of Kantorovich Potentials

### 31 Mar 2026 v1

@@ Abstract.
On the quantitative side, we control the diameter of the set of optimal dual potentials by the Hausdorff distance between the support of one of the marginal measures and a regular connected set in $\mathbb{R}^{d}$. In this way, we quantify the extent to which optimisers are almost unique when the support of one marginal measure is almost connected. This is a consequence of a novel characterisation of the set of optimal dual potentials as the intersection of an explicit family of half-spaces.
Key words and phrases:
optimal transport, Kantorovich potentials, duality, uniqueness
2020 Mathematics Subject Classification:
49Q22, 49K40
@@ 1.1. Contributions to qualitative uniqueness
Let $\rho$ and $\mu$ be probability measures on $\mathbb{R}^{d}$ such that $\operatorname{spt}\rho$ is rectifiably connected and $\operatorname{spt}\mu$ is bounded. Suppose that $c\in\mathcal{C}^{1}_{\operatorname{loc}}(\mathbb{R}^{d}\times\mathbb{R}^{d})$ and that there exist real-valued upper semi-continuous functions $a\in L^{1}(\rho)$ and $b\in L^{1}(\mu)$ such that
@@
Let $c\in C^{1}_{\operatorname{loc}}(\mathbb{R}^{d}\times\mathbb{R}^{d})$. Let $\rho$ and $\mu$ be probability measures on $\mathbb{R}^{d}$ with bounded support. Assume that $(A_{i})_{i\in I}$ and $(B_{j})_{j\in J}$ are families of measurable, rectifiably connected sets such that
@@ 1.2. Contributions to quantitative uniqueness
Our main quantitative uniqueness result is Theorem 1.4 below, which bounds $\operatorname{diam}_{L^{\infty}}\Phi_{c}(\rho,\mu)$ by the Hausdorff distance between $\operatorname{spt}\rho$ and a connected set $\Omega\subseteq\mathbb{R}^{d}$ satisfying Assumption 5.1. Roughly speaking, such sets are path-connected, using arcs of uniformly bounded length and curvature. Examples of sets verifying Assumption 5.1 include bounded $\mathcal{C}^{0}$ domains (domains whose boundary is locally the graph of a continuous function) and $\mathcal{C}^{1,1}$ compact connected submanifolds with boundary, see Section 5.
@@
In particular, when $\mathcal{X},\mathcal{Y}\subseteq\mathbb{R}^{d}$ are both compact, Theorem 1.4 applies to the power cost $c(x,y)=\|x-y\|^{p}$ for all $p\in[2,+\infty)$ with $\alpha=1$, giving an exponent $1/2$ for $d_{H}(\operatorname{spt}\rho,\Omega)$ at small scales. When $p\in(1,2]$, Theorem 1.4 applies with $\alpha=p-1$ and exponent 11/p1-1/p for $d_{H}(\operatorname{spt}\rho,\Omega)$, since the function $z\mapsto\|z\|^{p}$ is $\mathcal{C}^{1,p-1}_{\operatorname{loc}}$. (We provide a short proof due to [mischler2024quantitative] of this fact in Appendix A.) We do not know if these exponents are optimal. In Section 5.4, we provide an example showing that, in general, we cannot expect a better exponent than $1$.
@@
  • Absolutely continuous $\rho$, with density uniformly bounded below with respect to the uniform measure on a bounded Lipschitz domain. Indeed, bounded Lipschitz domains satisfy Assumption 5.1 by Proposition 5.2, and $\rho$ satisfies (1.8) with $s=d$, since the restriction of the Lebesgue measure to bounded Lipschitz domains satisfies (1.8).
@@ 1.6. Structure of the paper
  • In Section 5, we prove the quantitative theorems, Theorems 1.4, 1.5, 1.6 and 1.7. We also provide an example which demonstrates that, in general, we cannot have quantitative uniqueness with an exponent better than $1$.
@@ 3. Qualitative uniqueness
3.1. Proof of Theorem 1.2: Qualitative uniqueness with $\mathcal{C}^{1}_{\operatorname{loc}}$ cost
@@
Let $\rho$ and $\mu$ be probability measures on $\mathbb{R}^{d}$, with $\operatorname{spt}\mu$ bounded. Suppose that $c\in\mathcal{C}^{1}_{\operatorname{loc}}(\mathbb{R}^{d}\times\mathbb{R}^{d})$ and
@@
That $c\in\mathcal{C}^{1}_{\operatorname{loc}}$ satisfying (3.1) satisfies Assumption 2.1 is direct from our hypotheses. Since we assume that $\operatorname{spt}\mu$ is bounded, for any $K\subseteq\mathbb{R}^{d}$ compact we have $\|\nabla_{x}c\|_{\mathcal{C}^{0}(K\times\operatorname{spt}\mu)}<+\infty$ so that Assumption 3.1 holds also. The local Lipschitz property of potentials follows from standard envelope arguments; see [santambrogio2015optimal, Box 1.8]. It is not hard to show that if $\psi$ is optimal, then $\psi^{\overline{c},\operatorname{spt}\mu}\in L^{1}(\rho)$, see Appendix B. Thus, we cannot have $\psi^{\overline{c},\operatorname{spt}\mu}\equiv$-$\infty$. Since $\psi^{\overline{c},\operatorname{spt}\mu}$ is finite at at least one point, and is locally Lipschitz on $\mathbb{R}^{d}$, it is finite everywhere. ∎
@@
We use the hypothesis that $\operatorname{spt}\mu$ is bounded on two separate occasions, both corresponding to essentially the same phenomena. The first occasion is to ensure that for $c\in\mathcal{C}^{1}_{\operatorname{loc}}$, Assumption 3.1 holds, guaranteeing local Lipschitz regularity of $\rho$-side Kantorovich potentials. The second occasion is to guarantee that $\operatorname{proj}_{x}(\operatorname{spt}\gamma)=\operatorname{spt}\rho$ for any $\gamma\in\Pi(\rho,\mu)$.
@@ 4.2. Diameter bounds
Here $BV([0,1];\mathbb{R}^{d})$ denotes the space of all functions $u:[0,1]\to\mathbb{R}^{d}$ whose derivative is a finite, vector-valued Radon measure, see [evans2025measure, Chapter 5]. For $u\in BV([0,1];\mathbb{R}^{d})$ with vector measure derivative $\dot{u}$, we denote by $|\dot{u}|$ its total variation measure. We can always assume that we are dealing with a right continuous representative of $u$, so that for any $0\leq s<t\leq 1$, we have
$u(t)-u(s)=\dot{u}((s,t]).$
We have $BV([0,1];\mathbb{R}^{d})\subset L^{\infty}([0,1];\mathbb{R}^{d})$ and hence for $\omega:[0,1]\to\mathbb{R}^{d}$ with $\dot{\omega}\in BV([0,1];\mathbb{R}^{d})$, we automatically have $\omega\in W^{1,\infty}([0,1];\mathbb{R}^{d})$.
@@
which gives (4.12). Finally, we treat (4.17). For any $y\in\mathcal{Y}$ and $t\in[1/n,1]$, we have
@@ 5. Quantitative uniqueness
Motivated by the calculations of Lemma 4.4, we are interested in domains satisfying the following assumption.
Assumption 5.1 (Finite length and curvature path connectivity).
Suppose $\Omega\subseteq\mathbb{R}^{d}$ is such that
Sets satisfying Assumption 5.1 are path-connected, using paths of uniformly bounded length and curvature. The following classes of sets satisfy Assumption 5.1. We leave the proof to Appendix C.
Proposition 5.2.
@@
Let $\Omega=I_{1}\times\cdots I_{d}\subseteq\mathbb{R}^{d}$ for (potentially unbounded) intervals $I_{1},...,I_{d}\subseteq\mathbb{R}$. Then $\Omega$ satisfies Assumption 5.3 with $N(\Omega)=d$ and $S(\Omega)=1$,
@@ 5.4. Example: Optimality of the exponent $1$ for the grid case
then for each $n\in\mathbb{N}$, $\operatorname{spt}\rho_{n}=\mathcal{G}_{n}$. Denote the $x_{d}$-face of the cube $0,1]^{d}$ by $F=[0,1]^{d-1}\times\{1\}$. Then the primal quadratic optimal transport between $\rho_{n}$ to $\mu_{n}$ has a unique solution, which transports mass via

### 02 Apr 2026 v2

@@ Abstract.
On the quantitative side, we control the diameter of the set of optimal dual potentials by the Hausdorff distance between the support of one of the marginal measures and a regular connected set. In this way, we quantify the extent to which optimisers are almost unique when the support of one marginal measure is almost connected. This is a consequence of a novel characterisation of the set of optimal dual potentials as the intersection of an explicit family of half-spaces.
Keywords: optimal transport, Kantorovich potentials, duality, uniqueness
2020 Mathematics Subject Classification: 49Q22, 49K40
1CMAP École Polytechnique, Palaiseau, France. Email: william.ford@polytechnique.edu
@@ 1.1. Contributions to qualitative uniqueness
Let $\rho$ and $\mu$ be probability measures on $\mathbb{R}^{d}$ such that $\operatorname{spt}\rho$ is rectifiably connected and $\operatorname{spt}\mu$ is bounded. Suppose that $c\in\mathcal{C}^{1}(\mathbb{R}^{d}\times\mathbb{R}^{d})$ and that there exist real-valued upper semi-continuous functions $a\in L^{1}(\rho)$ and $b\in L^{1}(\mu)$ such that
@@
Let $c\in\mathcal{C}^{1}(\mathbb{R}^{d}\times\mathbb{R}^{d})$. Let $\rho$ and $\mu$ be probability measures on $\mathbb{R}^{d}$ with bounded support. Assume that $(A_{i})_{i\in I}$ and $(B_{j})_{j\in J}$ are families of measurable, rectifiably connected sets such that
@@ 1.2. Contributions to quantitative uniqueness
Our main quantitative uniqueness result bounds $\operatorname{diam}_{L^{\infty}}\Phi_{c}(\rho,\mu)$ by the Hausdorff distance between $\operatorname{spt}\rho$ and a connected set $\Omega\subseteq\mathbb{R}^{d}$ satisfying the following assumption.
Assumption 1.4.
Suppose $\Omega\subseteq\mathbb{R}^{d}$ is such that
Here $BV([0,1];\mathbb{R}^{d})$ denotes the space of functions whose derivative is a finite, vector-valued Radon measure, and $|\ddot{\omega}|$ denotes the variation measure of $\ddot{\omega}$, see [evans2025measure, Chapter 5]. Roughly speaking, such sets are rectifiably connected, using arcs of uniformly bounded length and curvature. Examples include bounded $\mathcal{C}^{0}$ domains (domains whose boundary is locally the graph of a continuous function) and $\mathcal{C}^{1,1}$ compact connected submanifolds with boundary, see Appendix C.
@@
In particular, when $\mathcal{X},\mathcal{Y}\subseteq\mathbb{R}^{d}$ are both compact, Theorem 1.5 applies to the cost $c(x,y)=\|x-y\|^{p}$ for all $p\in[2,+\infty)$ with $\alpha=1$, giving an exponent $1/2$ for $d_{H}(\operatorname{spt}\rho,\Omega)$ at small scales. When $p\in(1,2]$, Theorem 1.5 applies with $\alpha=p-1$ and exponent 11/p1-1/p for $d_{H}(\operatorname{spt}\rho,\Omega)$, since the function $z\mapsto\|z\|^{p}$ is $\mathcal{C}^{1,p-1}_{\operatorname{loc}}$. (We provide a short proof due to [mischler2024quantitative] of this fact in Appendix A.) We do not know if these exponents are optimal. In Section 5.4, we provide an example showing that, in general, we cannot expect a better exponent than $1$.
@@
  • Absolutely continuous $\rho$, with density uniformly bounded below with respect to the uniform measure on a bounded Lipschitz domain. Indeed, bounded Lipschitz domains satisfy Assumption 1.4 by Proposition C.1, and $\rho$ satisfies (1.9) with $s=d$, since the restriction of the Lebesgue measure to bounded Lipschitz domains satisfies (1.9).
@@ 1.6. Structure of the paper
  • In Section 5, we prove the quantitative theorems, Theorems 1.5, 1.6, 1.7 and 1.8. We also provide an example which demonstrates that, in general, we cannot have quantitative uniqueness with an exponent better than $1$.
@@ 3. Qualitative uniqueness
3.1. Proof of Theorem 1.2: Qualitative uniqueness with $\mathcal{C}^{1}$ cost
@@
Let $\rho$ and $\mu$ be probability measures on $\mathbb{R}^{d}$, with $\operatorname{spt}\mu$ bounded. Suppose that $c\in\mathcal{C}^{1}(\mathbb{R}^{d}\times\mathbb{R}^{d})$ and
@@
That $c\in\mathcal{C}^{1}$ satisfying (3.1) satisfies Assumption 2.1 is direct from our hypotheses. Since we assume that $\operatorname{spt}\mu$ is bounded, for any $K\subseteq\mathbb{R}^{d}$ compact we have $\|\nabla_{x}c\|_{\mathcal{C}^{0}(K\times\operatorname{spt}\mu)}<+\infty$ so that Assumption 3.1 holds also. The local Lipschitz property of potentials follows from standard envelope arguments; see [santambrogio2015optimal, Box 1.8]. It is not hard to show that if $\psi$ is optimal, then $\psi^{\overline{c},\operatorname{spt}\mu}\in L^{1}(\rho)$, see Appendix B. Thus, we cannot have $\psi^{\overline{c},\operatorname{spt}\mu}\equiv$-$\infty$. Since $\psi^{\overline{c},\operatorname{spt}\mu}$ is finite at at least one point, and is locally Lipschitz on $\mathbb{R}^{d}$, it is finite everywhere. ∎
@@
We use the hypothesis that $\operatorname{spt}\mu$ is bounded on two separate occasions, both corresponding to essentially the same phenomena. The first occasion is to ensure that for $c\in\mathcal{C}^{1}$, Assumption 3.1 holds, guaranteeing local Lipschitz regularity of $\rho$-side Kantorovich potentials. The second occasion is to guarantee that $\operatorname{proj}_{x}(\operatorname{spt}\gamma)=\operatorname{spt}\rho$ for any $\gamma\in\Pi(\rho,\mu)$.
@@ 4.2. Diameter bounds
which gives (4.12). (Here we used that without loss of generality we can assume we are dealing with a right continuous representative of $\dot{\omega}$, so that for any $0\leq s<t\leq 1$ we have $\dot{\omega}(t)-\dot{\omega}(s)=\ddot{\omega}((s,t])$.) Finally, we treat (4.17). For any $y\in\mathcal{Y}$ and $t\in[1/n,1]$, we have
@@ 5. Quantitative uniqueness
Let $\Omega=I_{1}\times\cdots I_{d}\subseteq\mathbb{R}^{d}$ for (potentially unbounded) intervals $I_{1},...,I_{d}\subseteq\mathbb{R}$. Then $\Omega$ satisfies Assumption 5.1 with $N(\Omega)=d$ and $S(\Omega)=1$.
@@ 5.4. Example: Optimality of the exponent $1$ for the grid case
then for each $n\in\mathbb{N}$, $\operatorname{spt}\rho_{n}=\mathcal{G}_{n}$. Denote the $x_{d}$-face of the cube $0,1]^{d}$ by $F=[0,1]^{d-1}\times\{1\}$. Then the primal quadratic optimal transport between $\rho_{n}$ and $\mu_{n}$ has a unique solution, which transports mass via